# Enhancement of non-Gaussianity and nonclassicality of pair coherent states with postselected von Neumann measurement

### v1

@@ VI Joint Wigner function
where $\rho$ is the density operator of the given state, and $P$ is the parity operator. However, in our current work, this Wigner function cannot be used directly since the PCS is a two-mode radiation field. In the two-mode case, the Wigner function of our state $|\Psi\rangle$ can be expressed in terms of the joint Wigner function
@@
where $\rho=|\Psi\rangle\langle\Psi|$ is the density operator of the state $|\Psi\rangle$, $P_{j}=P_{a}P_{b}=e^{i\pi a^{\dagger}a}e^{i\pi b^{\dagger}b}$ is the joint photon number parity operator, and $D(\alpha)=e^{\alpha a^{\dagger}-\alpha^{\ast}a}$ and $D(\beta)=e^{\beta b^{\dagger}-\beta^{\ast}b}$ are the displacement operators acting on the $a$ and $b$ modes of the two-mode state $|\Psi\rangle$, respectively. The joint Wigner function $W_{J}$ is a function in the $4D$ phase space, whose coordinates are ($Re(\alpha)$, $Im(\alpha)$, $Re(\beta)$, $Im(\beta)$). In Eq. (36), $P_{J}(\alpha,\beta)$ is referred to as the scaled Wigner function, which we use for our analysis of the phase space distribution of the measurement output state $|\Psi\rangle$. Upon examining the expression for $P_{J}(a,\beta)$, we find that it is the average value of the joint parity operator $P_{j}$ after applying displacement operators with amplitudes $-\alpha$ and $-\beta$ to the $a$ and $b$ modes, respectively. Since the eigenstates of the parity operators $P_{a}$ and $P_{b}$ are the Fock states $|n\rangle_{a}$ and $|n\rangle_{b}$, with eigenvalues $(-1)^{n_{a}}$ and $(-1)^{n_{b}}$, the scaled Wigner function is therefore bounded by ±$\pm 1$, i.e.,
@@
Figure 10: Joint Wigner function of the enhanced PCS state $|\Psi\rangle$ after postselected measurement.(a and b) A 2D plane cut along (a) the $Re(\alpha)-Re(\beta)$ axes and (b) the $Im(\alpha)-Im(\beta)$ axes of the calculated $4D$ scaled joint Wigner function $P_{J}(\alpha,\beta)$ for the measurement output state $|\Psi\rangle$ with $\Gamma=0.3$, $\gamma=0.5$, $\langle\sigma_{x}\rangle_{w}=5.761$ and $\delta=2$. (c) Diagonal line cuts of the data shown in (a), corresponding to a $1D$ plot of the calculated scaled joint Wigner function $P_{J}(\alpha,\beta)$ along $Re(\alpha)=Re(\beta)$ with $Im(\alpha)=Im(\beta)=0$. (d) Diagonal line cuts of the data shown in (b), corresponding to a $1D$ plot of the calculated scaled joint Wigner function $P_{J}(\alpha,\beta)$ along $Im(\alpha)=Im(\beta)$ with $Re(\alpha)=Re(\beta)=0$.
@@
To illustrate the core features in this 4D Wigner function of the state $|\Psi\rangle$ and compare it with the initial PCS state $|\phi\rangle$, we first present the 2D cuts along the $Re(\alpha)-Re(\beta)$ plane and the $Im(\alpha)-Im(\beta)$ plane of the joint Wigner function of $|\phi\rangle$, as shown in Fig. 9. Figs. 9(a) and 9(b) represent the density plot of the joint Wigner function of the initial PCS state $|\phi\rangle$ with $\gamma=0.5$ and $\delta=0$, in the planes $Im(\alpha)=Im(\beta)=0$ and $Re(\alpha)=Re(\beta)=0$, respectively. As shown in these figures, the initial PCS $|\phi\rangle$ is Gaussian when there is no PND between the two modes ($\delta=0$), and it exhibits squeezing along the $F_{2}$ quadrature (see Sec.III A). Similarly, Figs. 9(c) and 9(d) show the density plots of the joint Wigner function of the initial PCS state $|\phi\rangle$ with $\gamma=0.5$ and $\delta=2$, in the planes $Im(\alpha)=Im(\beta)=0$ and $Re(\alpha)=Re(\beta)=0$, respectively. It is evident that when there is PND between the two modes ($\delta=2$), the state $|\phi\rangle$ becomes non-Gaussian. Additionally, the characteristic interference fringes of the PCS are visible, indicating substantial coherence and consistent multiphoton phases. This result suggests that PND ($\delta\neq 0$) between the two modes is crucial for enhancing the nonclassicality of the PCS state $|\phi\rangle$.
@@
In Fig. 10, we present the scaled joint Wigner function $P_{J}(\alpha,\beta)$ of the state $|\Psi\rangle$ with $\Gamma=0.3$, $\gamma=0.5$, $\langle\sigma_{x}\rangle_{w}=5.761$, and $\delta=2$. We observe that this Wigner function exhibits highly nonclassical characteristics in phase space. Compared to the initial PCS state $|\phi\rangle$ (see Figs. 9(c) and 9(d)), we can see that after the postselected measurement, the shapes of the scaled joint Wigner functions not only show good squeezing for the appropriate coupling strength parameter $\Gamma$ and large anomalous weak value $\langle\sigma_{x}\rangle_{w}$, but we also can observe clear quantum interference structures formed between the peaks (see Figs. 10(c) and 10(d)). Furthermore, if someone take comparison Fig. 9 and Fig. 10 who can find that after postselected von Neumann measurement the negative values of Wigner function increased for large anomalous weak value compared to initial state case. This indicated that the nonclassicality of the PCS enhanced after the postselected von Neumann measurement.
@@
These phase space analyses demonstrate that after the postselected von Neumann measurement, the nonclassicality and non-Gaussianity of our PCS state $|\phi\rangle$ enhanced due to the weak signal amplification feature of the weak value.
@@ VII Fidelity
In this section, we examine the state distance between the initial PCS and final MD states, specifically $|\phi\rangle$ and $|\Psi\rangle$ using the fidelity function. The fidelity $F$ is defined as the square of the absolute value of their scalar product, and it is bounded as $0\leq F\leq 1$, where the lower (higher) boundary represents completely different (identical) states. We define the fidelity function between $|\phi\rangle$ and $|\Psi\rangle$ as:
@@ VIII Discussion
From the fidelity analysis of our measurement output, the enhanced state $|\Psi\rangle$, we can deduce that after the postselected von Neumann measurement, the initial PCS $|\phi\rangle$ undergoes a dramatic change, potentially transforming into a completely different state. As previous related studies have shown, the postselected von Neumann measurement technique provides an alternative approach to state optimization problems (Hu et al., 2022). In this context, we explore the concrete changes and potential applications of our enhanced state $|\Psi\rangle$ in state preparation processes. Depending on the coupling strength parameter $\Gamma$ and the state parameter $\gamma$, we outline its usefulness in the following four scenarios:
(1) Arbitrary values for the state parameter $\gamma$ and the coupling strength parameter $\Gamma$. In this general case, our final state is the measurement output state $|\Psi\rangle$ as defined in Eq. (10). For simplicity, the initial PCS $|\phi\rangle$ expressed in the coherent state representation as (Agarwal, 1988).
$|\phi\rangle=\mathcal{N}_{n}e^{{}^{|\gamma|}}\int\frac{\mathrm{d}\theta}{2\pi}\left[(\gamma)^{1/2}e^{i\theta}\right]^{-\delta}|(\gamma)^{1/2}e^{i\theta}\rangle_{a}|(\gamma)^{1/2}e^{-i\theta}\rangle_{b}.$ (39)
@@
$|\Psi_{1}\rangle$ $=\frac{\lambda}{2}\left[t_{+}|\gamma+\frac{\Gamma_{a}}{2},\delta\rangle+t_{-}|\gamma-\frac{\Gamma_{a}}{2},\delta\rangle\right].$ (40)
Here, $t_{\pm}=1\pm\langle\sigma_{x}\rangle_{w}$ and the state $|\gamma\pm\frac{\Gamma_{a}}{2},\delta\rangle$ represents the displaced PCS (the displacement operator $D(\pm\frac{\Gamma}{2})$ acts only on $a$ mode), and its expression is given by:
$|\gamma\pm\frac{\Gamma_{a}}{2},\delta\rangle$ $=\frac{\mathcal{N}_{n}e^{|\gamma|}}{2\pi}\int\left[(\gamma)^{1/2}e^{i\theta}\right]^{-\delta}e^{\pm\frac{\Gamma}{2}[iIm[((\gamma)^{1/2}e^{-i\theta})^{*}]]}$ (41)
@@
(2) Strong coupling strength parameter ($\Gamma\gg 1$) with small parameter $\gamma$ $(|\gamma|\ll 1)$ Since the state parameter $\gamma$ is very small, we expand the Eq. (1) up to the first order of $\gamma$ and ignore higher order terms. In this extreme case, the Eq. (10) changed to:
@@
$|\Psi^{{}^{\prime}}\rangle$ $\thickapprox\frac{\lambda^{{}^{\prime}}}{2}\left\{\left[t_{+}^{{}^{\prime}}\left|\frac{\Gamma}{2},\delta\right\rangle_{a}+t_{-}^{{}^{\prime}}\left|-\frac{\Gamma}{2},\delta\right\rangle_{a}\right]\left|0\right\rangle_{b}\right.$
$+\left.\frac{\gamma}{\sqrt{1+\delta}}\left[t_{+}^{{}^{\prime}}\left|\frac{\Gamma}{2},1+\delta\right\rangle_{a}+t_{-}^{{}^{\prime}}\left|-\frac{\Gamma}{2},1+\delta\right\rangle_{a}\right]\left|1\right\rangle_{b}\right\},$ (42)
@@
where $\lambda^{{}^{\prime}}=\frac{\sqrt{2}\gamma^{\delta/2}}{\sqrt{\delta!}}\left[1+|\langle\sigma_{x}\rangle_{c}|^{2}+(1-|\langle\sigma_{x}\rangle_{c}|^{2})P\right]^{-\frac{1}{2}}$ and $t_{\pm}^{{}^{\prime}}=1\pm\langle\sigma_{x}\rangle_{c}$. The states |±$\left|\pm\frac{\Gamma}{2},\delta\right\rangle_{a}=D(\pm\frac{\Gamma}{2})|\delta\rangle_{a}$ and |±$\left|\pm\frac{\Gamma}{2},1+\delta\right\rangle_{a}=D(\pm\frac{\Gamma}{2})|1+\delta\rangle_{a}$ denote the displaced $\delta$ and $1+\delta$ photon Fock states of the $a$-mode, respectively. Here, $\langle\sigma_{x}\rangle_{c}$ represents the conditional expectation value of the measured system observable $\sigma_{x}$, which is measured through the postselected strong measurement. For our pre- and post-selected system states $|\psi_{i}\rangle$ and $|\psi_{f}\rangle$, as defined in Sec. II, this conditional expectation value is given by $\langle\sigma_{x}\rangle_{c}=\cos\vartheta\sin\alpha$ (eq, ). It can be observed that the state $|\Psi^{{}^{\prime}}\rangle$ is entangled state of the $a$ and $b$-modes, and interesting quantum phenomena in the $a$-mode can be explored by adjusting the related parameters.
@@
If the expectation value $\langle\sigma_{x}\rangle_{c}=0$ , i.e. $\alpha=0$, $\vartheta=\frac{\pi}{2}$ or $\frac{3\pi}{2},$ the Eq. (42) changed to $|\Psi^{\prime}\rangle\thickapprox\left(\left|\frac{\Gamma}{2},\delta\right\rangle_{a}+\left|-\frac{\Gamma}{2},\delta\right\rangle_{a}\right)\left|0\right\rangle_{b}$, with high probability due to the small state parameter $\gamma$. In this case, the $b$-mode is in the vacuum state, and the $a$-mode is in superposition of displaced Fock state, i.e., $\left|\frac{\Gamma}{2},\delta\right\rangle_{a}+\left|-\frac{\Gamma}{2},\delta\right\rangle_{a}$. One interesting point is that if the PND between the two modes is zero ($\delta=0$), the $a$ mode is prepared in even Schrödinger cat state $\left|\frac{\Gamma}{2}\right\rangle_{a}+\left|-\frac{\Gamma}{2}\right\rangle_{a}.$ Furthermore, if $\langle\sigma_{x}\rangle_{c}=1$, the above state became another entangled state as $|\frac{\Gamma}{2},\delta\rangle_{a}|0\rangle_{b}+\frac{\gamma}{\sqrt{1+\delta}}|\frac{\Gamma}{2},1+\delta\rangle_{a}|1\rangle_{b}.$ If we further assume $|\gamma|\ll 1$, then $a$ mode stated in displaced Fock state $D(\frac{\Gamma}{2})|\delta\rangle_{a}$ meanwhile the $b$ mode in the vacuum state.
@@
(3) Weak coupling strength parameter ($\Gamma\ll 1$) with state parameter $\gamma$ ($|\gamma|\gg 1$). In this case, since the coupling strength $\Gamma$ is small, we can perform a first-order Taylor expansion of the associated operators; The displacement operator is approximately $D(\frac{\Gamma}{2})\approx\mathbb{I}+\frac{\Gamma}{2}\left(a^{\dagger}-a\right)$. Then, the final MD state $|\Psi\rangle$ (Eq. (10)) changes to (unnormalized):

### v2

@@ VI Joint Wigner function
where $\rho$ is the density operator of the given state, and $P=e^{i\pi a^{\dagger}a}$ is the parity operator. However, in our current work, this Wigner function cannot be used directly since the PCS is a two-mode radiation field. In the two-mode case, the Wigner function of the final MD state $|\Psi\rangle$ can be expressed in terms of the joint Wigner function
@@
where $\rho=|\Psi\rangle\langle\Psi|$ is the density operator of the state $|\Psi\rangle$, $P_{j}=P_{a}P_{b}=e^{i\pi a^{\dagger}a}e^{i\pi b^{\dagger}b}$ is the joint photon number parity operator, and $D(\alpha)=e^{\alpha a^{\dagger}-\alpha^{\ast}a}$ and $D(\beta)=e^{\beta b^{\dagger}-\beta^{\ast}b}$ are the displacement operators acting on the $a$ and $b$ modes of the two-mode state $|\Psi\rangle$, respectively. The joint Wigner function $W_{J}$ is a function in the $4D$ phase space, whose coordinates are ($Re[\alpha]$, $Im[\alpha]$, $Re[\beta]$, $Im[\beta]$). In Eq. (36), $P_{J}(\alpha,\beta)$ is referred to as the scaled Wigner function, which we use for our analysis of the phase space distribution of the measurement output state $|\Psi\rangle$. Upon examining the expression for $P_{J}(a,\beta)$, we find that it is the average value of the joint parity operator $P_{j}$ after applying displacement operators with amplitudes $-\alpha$ and $-\beta$ to the $a$ and $b$ modes, respectively. Since the eigenstates of the parity operators $P_{a}$ and $P_{b}$ are the Fock states $|n\rangle_{a}$ and $|n\rangle_{b}$, with eigenvalues $(-1)^{n_{a}}$ and $(-1)^{n_{b}}$, the scaled Wigner function is therefore bounded by ±$\pm 1$, i.e.,
@@
Figure 10: Joint Wigner function of the final MD state $|\Psi\rangle$ after postselected measurement.(a and b) A 2D plane cut along (a) the $Re[\alpha]-Re[\beta]$ axes and (b) the $Im[\alpha]-Im[\beta]$ axes of the calculated $4D$ scaled joint Wigner function $P_{J}(\alpha,\beta)$ for the measurement output state $|\Psi\rangle$ with $\Gamma=0.3$, $\gamma=0.5$, $\langle\sigma_{x}\rangle_{w}=5.761$ and $\delta=2$. (c) Diagonal line cuts of the data shown in (a), corresponding to a $1D$ plot of the calculated scaled joint Wigner function $P_{J}(\alpha,\beta)$ along $Re[\alpha]=Re[\beta]$ with $Im[\alpha]=Im[\beta]=0$. (d) Diagonal line cuts of the data shown in (b), corresponding to a $1D$ plot of the calculated scaled joint Wigner function $P_{J}(\alpha,\beta)$ along $Im[\alpha]=Im[\beta]$ with $Re[\alpha]=Re[\beta]=0$.
@@
To illustrate the core features in this 4D Wigner function of the final MD state $|\Psi\rangle$ and compare it with the initial PCS state $|\phi\rangle$, we first present the 2D cuts along the $Re[\alpha]-Re[\beta]$ plane and the $Im[\alpha]-Im[\beta]$ plane of the joint Wigner function of $|\phi\rangle$, as shown in Fig. 9. Figs. 9(a) and 9(b) represent the density plot of the joint Wigner function of the initial PCS state $|\phi\rangle$ with $\gamma=0.5$ and $\delta=0$, in the planes $Im[\alpha]=Im[\beta]=0$ and $Re[\alpha]=Re[\beta]=0$, respectively. As shown in these figures, the initial PCS $|\phi\rangle$ is Gaussian when there is no PND between the two modes ($\delta=0$), and it exhibits squeezing along the $F_{2}$ quadrature [see Sec.III A]. Similarly, Figs. 9(c) and 9(d) show the density plots of the joint Wigner function of the initial PCS state $|\phi\rangle$ with $\gamma=0.5$ and $\delta=2$, in the planes $Im[\alpha]=Im[\beta]=0$ and $Re[\alpha]=Re[\beta]=0$, respectively. It is evident that when there is PND between the two modes ($\delta=2$), the state $|\phi\rangle$ becomes non-Gaussian. Additionally, the characteristic interference fringes of the PCS are visible, indicating substantial coherence and consistent multiphoton phases. This result suggests that PND ($\delta\neq 0$) between the two modes is crucial for enhancing the nonclassicality of the PCS state $|\phi\rangle$.
@@
In Fig. 10, we present the scaled joint Wigner function $P_{J}(\alpha,\beta)$ of the final MD state $|\Psi\rangle$ with $\Gamma=0.3$, $\gamma=0.5$, $\langle\sigma_{x}\rangle_{w}=5.761$, and $\delta=2$. We observe that this Wigner function exhibits highly nonclassical characteristics in phase space. Compared to the initial PCS state $|\phi\rangle$ [see Figs. 9 (c) and 9 (d)], we can see that after the postselected measurement, the shapes of the scaled joint Wigner functions not only show good squeezing for the appropriate coupling strength parameter $\Gamma$ and large anomalous weak value $\langle\sigma_{x}\rangle_{w}$, but we also can observe clear quantum interference structures formed between the peaks [see Figs. 10 (c) and 10 (d)]. Furthermore, if someone take comparison Fig. 9 and Fig. 10 who can find that after postselected von Neumann measurement the negative values of Wigner function increased for large anomalous weak value compared to initial state case. This indicated that the nonclassicality of the PCS enhanced after the postselected von Neumann measurement.
@@
These phase space analyses demonstrate that after the postselected von Neumann measurement, the nonclassicality and non-Gaussianity of PCS state $|\phi\rangle$ is enhanced due to the weak signal amplification feature of the weak value.
@@ VII Fidelity
To deeply understand the effects of postselected von Neumann measurement on the initial PCS, in this section we study state distance between initial and final MD state corresponding to the before and after measurement process, and the teleportation fidelity.
VII.1 State distance
In this sub-section, we examine the state distance between the initial PCS and final MD states, specifically $|\phi\rangle$ and $|\Psi\rangle$ using the fidelity function. The fidelity $F$ is defined as the square of the absolute value of their scalar product, and it is bounded as $0\leq F\leq 1$, where the lower (higher) boundary represents completely different (identical) states. We define the fidelity function between $|\phi\rangle$ and $|\Psi\rangle$ as:
@@ VII.1 State distance
VII.2 Fidelity of teleportation
As discussed in Sec. V, the postselected von Neumann measurement can enhance the entanglement between two modes of PCS for some parameter regions. We know that the distillation or purification of entanglement are vital to realize the long-standing quantum communication. In order to check the effects of our von Neumann-type measurment on the efficiency of teleportation by using the PCS as the entangled channel, we use the fidelity to measure the performance of entanglement.
Here we consider the protocol of quantum teleportation was proposed by Vaidman, Braunstein and Kimble (called VBK protocol) (Vaidman, 1994; Braunstein and Kimble, 1998; Furusawa et al., 1998). We take the final MD state $|\Psi\rangle_{ab}$ [see the Eq. (10)] as the teleportation channel to teleport a coherent state by using VBK protocol. Next, we briefly describe the teleportation processes. We assume that as the sender and recipient the Alice and Bob possess and could control the $a$ mode and $b$ mode of the state $|\Psi\rangle_{ab}$, respectively, and the there have a coherent state $|\alpha\rangle_{c}$ in $c$ mode to be teleported. Thus, initially the composite three modes state takes the form $|\Psi_{1}\rangle=|\Psi\rangle_{ab}|\alpha\rangle_{c}$. Firstly, Alice take a joint measurement of the orthogonal quadrature components of modes $a$ and $c$ characterized by (Hofmann et al., 2000)
$|\beta\rangle_{ac}=\frac{1}{\sqrt{\pi}}\sum_{k=0}^{\infty}D_{c}(\beta)|k\rangle_{a}|k\rangle_{c},$ (39)
where $D_{c}(\beta)$ represents the displacement operator acting on mode $c$. After the joint measurement the state $|\Psi_{1}\rangle$ changes to $|\Psi_{2}\rangle={}_{ac}\langle\beta|\Psi_{1}\rangle$. Then, Alice communicate with Bob with classical information channel and tell the information of the parameter $\beta$. Finally, Bob takes a displacement operation over his $b$ mode according to the received value of $\beta$ from Alice. After all operations are completed, the resulting state at Bob’s side is given as
$|\Psi_{b}\rangle$ $=\frac{\mathcal{N}_{\delta}\lambda}{2\sqrt{\pi}}\sum_{n,m=0}^{\infty}\frac{\gamma^{n+\delta/2}h_{m}}{\sqrt{n!(n+\delta)!}}$
×$\times\left(t_{+}J_{+,mn}+t_{-}J_{-,mn}\right)D_{b}\left(\beta\right)|n\rangle_{b},$ (40)
where, $h_{m}=\langle m|D_{c}\left(-\beta\right)|\alpha\rangle_{c}$ , $t_{\pm}=1\pm\langle\sigma_{x}\rangle_{w}$ and $J_{\pm,mn}=\langle m|D_{a}\left(\pm\frac{\Gamma}{2}\right)|n+\delta\rangle_{a}$.
Figure 12: The fidelity of teleportation of coherent state for different system parameters. The (a) and (b) present the fidelity of teleportation as a function of PCS state parameter $\gamma$ for different values of $\alpha$ The (c) shows the fidelity of teleportation as a function of weak value parameter $\alpha$ for different coupling strength parameter $\Gamma$ while fixed $\gamma=3$ and $\delta=0$. In (a) $\delta=0$ and (b) $\delta=1$, and $\Gamma=1$ in both cases, and other parameters are the same as Fig. 2.
To evaluate this teleportation process we can rely on the average fidelity defined by
$F_{av}=\int d^{2}\beta|\langle\alpha|\Psi_{b}\rangle|^{2}.$ (41)
The above integral have exact value and its explicit expression is listed in Appendix A. The teleportation process is considered successful when $0.5<F_{av}\leq 1$ and the higher average fidelity indicates the better quality of the teleportation process. In Fig. 12 we plot the average fidelity $F_{av}$ of the teleportation as a function of $\gamma$ for different weak value parameter $\alpha$. The black curves in Fig. 12 (a) and (b) corresponding to the $\delta=0$ and $\delta=1$ cases of initial PCS $|\phi\rangle$. It can observe that the average fidelity $F_{av}$ significantly dependent on the PND $\delta$. When the PND is set to $\delta=0$, its corresponding average teleportation fidelity is fully higher than the case of $\delta=1$ for whole parameter regimes. For $\delta=0$ case the maximum value $0.7589$ of $F_{av}$ achieved at near $\gamma\approx 1.22$.
As showed in Fig. 12 (a), for $\delta=0$ case in small state parameter $\gamma$ regions the postselected von Neumann measurement have no positive effect on the teleportation task if the weak value $\langle\sigma_{x}\rangle_{w}$ takes anomalous values. However, with increasing the parameter $\gamma$ the average fidelity $F_{av}$ also increases along with large weak values. For $\delta=0$ case, if $\gamma>0.9$ the average fidelity always larger than the minimum value ($i.e.,$$0.5$) of successful teleportation and this claim is valid for all $\Gamma\leq 1$ regimes [see the Fig. 12 (c)].
In Fig. 12 (b) we presented the average fidelity $F_{av}$ for $\delta=1$ case. It can be seen that in small state parameter ($\gamma$<1) regions the postselection operation could let the $F_{av}$ take larger value than the initial PCS whereas those values can’t guarantee the successful teleportation. However, in $\gamma>1$ regions of $\delta=1$, same as $\delta=0$ case, after the postselected von Neumann measurement the average fidelity still can’t beyond the initial PCS. One interesting point in $\delta=1$ case is that with increase the state parameter $\gamma$ and weak value the $F_{av}$ also increasing, and can exceed 0.5 for large anomalous weak values and large $\gamma$.
In a word, compared with the initial PCS state $|\phi\rangle$ the postselected von Neumann measurement ($\Gamma\neq 0$) can’t enhance the performance of successful teleportation of coherent state if we take the state $|\Psi\rangle_{ab}$ as quantum channel. But, for PND $\delta=0$ and $\delta=1$ cases, the postselected von Neumann measurement also could guarantee to keep the successful teleportation in weak measurement regimes with anomalous weak values.
@@ VIII Discussion
From the fidelity analysis of our measurement output, the final MD state $|\Psi\rangle$, we can deduce that after the postselected von Neumann measurement, the initial PCS $|\phi\rangle$ undergoes a dramatic change, potentially transforming into a completely different state. As previous related studies have shown, the postselected von Neumann measurement technique provides an alternative approach to state engineering processes and guarantee the high fidelity (Hu et al., 2022). In most usual state preparation and state optimization methods the photon addition and /or photon subtraction are necessary (Kitagawa et al., 2006; Parigi et al., 2007; Biswas and Agarwal, 2007; Truong et al., 2021; Chuong and Duc, 2023), but if one use the postselected von Neumann measurement method to state preparation processes who can prepare the photon added or photon subtracted quantum states without adding or subtracting photons (Hu et al., 2022). In this context, we explore the concrete changes of the final MD state $|\Psi\rangle$ in state preparation processes. Depending on the coupling strength parameter $\Gamma$ and the state parameter $\gamma$, we outline its usefulness in the following four scenarios:
(1) Arbitrary values for the state parameter $\gamma$ and the coupling strength parameter $\Gamma$. In this general case, our final MD state is the measurement output state $|\Psi\rangle$ as defined in Eq. (10). For simplicity, the initial PCS $|\phi\rangle$ expressed in the coherent state representation as (Agarwal, 1988).
$|\phi\rangle=\mathcal{N}_{\delta}e^{{}^{|\gamma|}}\int\frac{\mathrm{d}\theta}{2\pi}\left[(\gamma)^{1/2}e^{i\theta}\right]^{-\delta}|(\gamma)^{1/2}e^{i\theta}\rangle_{a}|(\gamma)^{1/2}e^{-i\theta}\rangle_{b}.$ (42)
@@
$|\Psi\rangle$ $=\frac{\lambda}{2}\left[t_{+}|\gamma+\frac{\Gamma_{a}}{2},\delta\rangle+t_{-}|\gamma-\frac{\Gamma_{a}}{2},\delta\rangle\right].$ (43)
Here, the state $|\gamma\pm\frac{\Gamma_{a}}{2},\delta\rangle$ represents the displaced PCS (the displacement operator $D(\pm\frac{\Gamma}{2})$ acts only on $a$ mode), and its expression is given by:
$|\gamma\pm\frac{\Gamma_{a}}{2},\delta\rangle$ $=\frac{\mathcal{N}_{\delta}e^{|\gamma|}}{2\pi}\int\left[(\gamma)^{1/2}e^{i\theta}\right]^{-\delta}e^{\pm\frac{\Gamma}{2}[iIm[((\gamma)^{1/2}e^{-i\theta})^{*}]]}$ (44)
@@
The final MD state $|\Psi\rangle$ is the superposition of $|\gamma+\frac{\Gamma_{a}}{2},\delta\rangle$ and $|\gamma-\frac{\Gamma_{a}}{2},\delta\rangle$. By taking the postselection probability of measured system (it equals to $|\langle\psi_{f}|\psi_{i}\rangle|^{2}$) into account, the success probability of postselection of our MD state is (Wu and Żukowski, 2012; Liu et al., 2022)
$P_{s}$ $=\langle\phi^{\prime}|\phi^{\prime}\rangle$
$=\frac{1}{2}\cos^{2}\frac{\alpha}{2}\left[1+|\langle\sigma_{x}\rangle_{w}|^{2}+(1-|\langle\sigma_{x}\rangle_{w}|^{2})P\right]$ (45)
where $|\phi^{\prime}\rangle=\langle\psi_{f}|\Psi_{evol}\rangle$ and the expression of $|\Psi_{evol}\rangle$ is given in Eq. (II). For $\delta=0$ case, the dependence of $P_{s}$ on the weak value $\langle\sigma_{x}\rangle_{w}$ and coupling strength parameter $\Gamma$ for fixed state parameter $\gamma$ is presented in Fig. 13. As showed Fig. 13, the success probability of postselection of our MD state is not very low especially in very weak coupling and small weak values regimes. With increase the coupling strength parameter $\Gamma$ the $P_{s}$ is decreasing for normal weak values of system observable. Whereas, if $\langle\sigma_{x}\rangle_{w}$ take anomalous values ($\frac{\pi}{2}$$<\alpha<\pi$), the output state detected probability $P_{s}$ increased with increasing the coupling strength $\Gamma$.
Figure 13: The success probability of postselection of the final MD state $|\Psi\rangle$ as a function weak value parameter $\alpha$ for different coupling strength parameter $\Gamma.$ Here, we take $\delta=0,$ $\gamma=2$ and other parameters are the same as Fig. 2.
@@
(2) Strong coupling strength parameter ($\Gamma>1$) with small state parameter $\gamma$ $(|\gamma|\ll 1)$ Since the state parameter $\gamma$ is very small, we expand the Eq. (1) up to the first order of $\gamma$ and ignore higher order terms. In this extreme case, the Eq. (43) changed to:
@@
$|\Psi^{{}^{\prime}}\rangle$ $\thickapprox\frac{\lambda^{{}^{\prime}}}{2}\left\{\left[t_{+}\left|\frac{\Gamma}{2},\delta\right\rangle_{a}+t_{-}\left|-\frac{\Gamma}{2},\delta\right\rangle_{a}\right]\left|0\right\rangle_{b}\right.$
$+\left.\frac{\gamma}{\sqrt{1+\delta}}\left[t_{+}\left|\frac{\Gamma}{2},1+\delta\right\rangle_{a}+t_{-}\left|-\frac{\Gamma}{2},1+\delta\right\rangle_{a}\right]\left|1\right\rangle_{b}\right\},$ (46)
@@
where $\lambda^{{}^{\prime}}=\frac{\sqrt{2}\gamma^{\delta/2}}{\sqrt{\delta!}}\left[1+|\langle\sigma_{x}\rangle_{w}|^{2}+(1-|\langle\sigma_{x}\rangle_{w}|^{2})P\right]^{-\frac{1}{2}}$. The states |±$\left|\pm\frac{\Gamma}{2},\delta\right\rangle_{a}=D(\pm\frac{\Gamma}{2})|\delta\rangle_{a}$ and |±$\left|\pm\frac{\Gamma}{2},1+\delta\right\rangle_{a}=D(\pm\frac{\Gamma}{2})|1+\delta\rangle_{a}$ denote the displaced $\delta$ and $1+\delta$ photon Fock states of the $a$-mode, respectively. It can be observed that the state $|\Psi^{{}^{\prime}}\rangle$ is entangled state of the $a$ and $b$-modes, and interesting quantum phenomena in the $a$-mode can be explored by adjusting the related parameters.
If $\langle\sigma_{x}\rangle_{w}=0$ , i.e. $\alpha=0$, the Eq. (46) changed to
$|\Psi_{ch}^{\prime}\rangle\approx$$\left\{\left[\left|\frac{\Gamma}{2},\delta\right\rangle_{a}+\left|-\frac{\Gamma}{2},\delta\right\rangle_{a}\right]\left|0\right\rangle_{b}\right.$
$+\left.\frac{\gamma}{\sqrt{1+\delta}}\left[\left|\frac{\Gamma}{2},1+\delta\right\rangle_{a}+\left|-\frac{\Gamma}{2},1+\delta\right\rangle_{a}\right]\left|1\right\rangle_{b}\right\}.$ (47)
@@
In this expression the entangled state between two modes can be seen very clearly. If the $b$-mode is in the vacuum state $\left|0\right\rangle_{b}$, and the $a$-mode is in superposition of displaced Fock state, i.e., $\left|\frac{\Gamma}{2},\delta\right\rangle_{a}+\left|-\frac{\Gamma}{2},\delta\right\rangle_{a}$. In this case, an interesting point is that if the PND between the two modes is zero ($\delta=0$), the $a$ mode is prepared in even Schrödinger cat state $\left|\frac{\Gamma}{2}\right\rangle+\left|-\frac{\Gamma}{2}\right\rangle.$ If we take measurement on mode $b$ and detect a single photon, the $a$ mode prepared in the superposition of displaced number states:
$|\phi\rangle_{a}\thicksim\langle 1|_{b}\left(|\Psi_{ch}^{\prime}\rangle\right)\thicksim\left[D\left(\frac{\Gamma}{2}\right)+D\left(-\frac{\Gamma}{2}\right)\right]\left|1+\delta\right\rangle_{a},$ (48)
Furthermore, if $\langle\sigma_{x}\rangle_{w}=1$, the above state became another entangled state as
$|\psi_{ch}\rangle=|\frac{\Gamma}{2},\delta\rangle_{a}|0\rangle_{b}+\frac{\gamma}{\sqrt{1+\delta}}|\frac{\Gamma}{2},1+\delta\rangle_{a}|1\rangle_{b}.$ (49)
It can be seen that if the $b$-mode in the vacuum state $|0\rangle_{b}$ then $a$ mode stated in displaced PND state $D(\frac{\Gamma}{2})|\delta\rangle_{a}$ Upon detecting a single photon at mode $b$, mode $a$ collapses to displaced $1+\delta$ photon state:
$|\phi\rangle_{a}\thicksim\langle 1|_{b}\left(|\psi_{ch}\rangle\right)\thicksim D\left(\frac{\Gamma}{2}\right)|1+\delta\rangle_{a}.$ (50)
@@
(3) Weak coupling strength parameter ($\Gamma\ll 1$) with state parameter $\gamma$ ($|\gamma|>1$). In this case, since the coupling strength $\Gamma$ is small, we can perform a first-order Taylor expansion of the associated operators; The displacement operator is approximately $D(\frac{\Gamma}{2})\approx\mathbb{I}+\frac{\Gamma}{2}\left(a^{\dagger}-a\right)$. Then, the final MD state $|\Psi\rangle$ (Eq. (10)) changes to (unnormalized):